Mathematics 242 – Principles of Analysis Definitions Quiz 4 – October 8, 2004

Your Name:

1. (20) What does it mean to say a real number s is the supremum (or least upper bound) of a set S.

2. (5) What is the supremum of the set S = {2 − 1/n : n ∈ N}.